OpenStudy (anonymous):

MEDAL WITH ME GIVEN!!! A triangle has squares on its three sides as shown below. What is the value of x? 3 4 5 7

OpenStudy (xaze):

Where's the picture of the triangle?

OpenStudy (anonymous):

its loading

OpenStudy (xaze):

Okay, you're going to want to use the formula a^2+b^2=c^2

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

5?

OpenStudy (xaze):

a = 3 b = 4

OpenStudy (xaze):

a^2+b^2=c^2 \[3^2+4^2=c\]\[9+16=c\]

OpenStudy (misty1212):

\[3+4=7\\ 3^2+4^2=25\]

OpenStudy (misty1212):

\[\sqrt{25}=5\]

OpenStudy (anonymous):

i said that in the beginning thank you @misty1212

OpenStudy (anonymous):

@Xaze and you too.

OpenStudy (misty1212):

you are welcome

OpenStudy (xaze):

Yup